A 32.0-Ω resistor and a 20.0-Ω are connected in parallel, and the combination is applied



across a 240 V dc line.



(a) What is the resistance of the parallel combination?



(b) What is the total current through the parallel combination?



(c) What is the current through each resistor?




1
Expert's answer
2022-05-04T12:40:29-0400

a) Resistance of the parallel combination:

"R = \cfrac{R1*R2}{R1+R2} = \cfrac{32.0*20.0}{32.0+20.0} \approx 12.3 (Ohm)"



b) Total current through the parallel combination

"I = \cfrac{V}{R} = \cfrac{240}{ 12.3} \approx 19.5 (A)"



c) Current through each resistor

"I1 = \cfrac{V}{R1} = \cfrac{240}{32.0} = 7.5 (A)\\

I2 = \cfrac{V}{R2} = \cfrac{240}{20.0} = 12 (A)"